OpenStudy (anonymous):

Help please A client comes to you for investment advice on his $500,000 winnings from the lottery. He has been offered the following options by three different financial institutions and requests assistance to help understand which option would be the best for his investment. Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years. Option 3: 14.5% simple interest for 10 years.

OpenStudy (anonymous):

At the end of 5 years: 1. 6/4 = 1.5 for 20 periods 1.015^20 = 1.35 gained 35% in 5 years 2. 1.08^5 = 1.48 gained 48% in 5 years 3. 14.5*5 = gained 72.5% in 5 years (best) Compounding is great but does not make up for this huge difference in interest rates.